I am having trouble with this question…Use the arithmetic series of numbering of 1,3,5,7,9,….to find the following: What is the difference between any 2 terms?
Is this the formula for this question….Sum= n(a1+an)/2) if this right I’m not sure where to go next. Please help…..

Answer by venugopalramana(3286)

I am having trouble with this question…Use the arithmetic series of numbering of 1,3,5,7,9,….to find the following: What is the difference between any 2 terms?
TAKE ANY 2 CONSECUTIVE TERMS AND SUBTRACT THE FORMER FROM LATER.THAT IS CALLED THE COMMON DIFFERENCE WHICH IS CONSTANT IN AN ARITHMATIC PROGRESSION.HERE WE HAVE
3-1=2 AND ALSO
5-3=7-5=9-7=2 HENCE THIS IS AN AP WITH CD OF 2 AND FIRST TERM AS 1.
Is this the formula for this question….Sum= n(a1+an)/2) if this right I’m not sure where to go next. Please help…..
THERE IS NO NEED FOR THIS FORMULA